OpenStudy (shaik0124):

The smallest number, which is a perfect square and contains 7936 as a factor is: Option 1 : 251664 Option 2 : 231564 Option 3 : 246016 Option 4 : 346016 #will give medal

OpenStudy (ikram002p):

i would try them all :D

OpenStudy (ikram002p):

with a tiny calculator

ganeshie8 (ganeshie8):

maybe start by finding the prime factorization of 7936

OpenStudy (dan815):

^

OpenStudy (ikram002p):

or simply check from calculator xD

OpenStudy (dan815):

2^8 * 31=7936

OpenStudy (dan815):

so 7936*31 = largest square

OpenStudy (dan815):

as u can split it into (2^4*31)(2^4*31)

OpenStudy (dan815):

i mean smallest square

OpenStudy (ikram002p):

or use long division

OpenStudy (ikram002p):

dan so wrong

OpenStudy (ikram002p):

the smallest is 2^4*(31^2)

OpenStudy (dan815):

oh true

OpenStudy (ikram002p):

=)

OpenStudy (dan815):

no its not

OpenStudy (dan815):

7396 is not a factor

ganeshie8 (ganeshie8):

7936 doesnt divide 2^4*(31^2)

OpenStudy (dan815):

sheeshh no one gets my sarcasm

OpenStudy (ikram002p):

:O i used dan factors

OpenStudy (dan815):

haha

OpenStudy (dan815):

2^8*31 must exist! in our square no question about it

OpenStudy (ikram002p):

-.-

OpenStudy (shaik0124):

3 is the answer

OpenStudy (dan815):

ya good

OpenStudy (shaik0124):

7936*31=246016

OpenStudy (ikram002p):

7936=2^8 *31 smallest perfect 2*8*31 *31

OpenStudy (shaik0124):

n a division problem, the divisor is twenty times the quotient and five times the remainder. If remainder is 16, the number will be: Option 1 : 3360 Option 2 : 336 Option 3 : 1616 Option 4 : 20516

ganeshie8 (ganeshie8):

one question per thread, please